(2+ycosx)dx+cosxdy=0


1
Expert's answer
2021-10-06T10:18:53-0400
"y'+y=-\dfrac{2}{\cos x}"

Integrating factor


"\mu(x)=e^x"

"e^xy'+e^xy=-\dfrac{2e^x}{\cos x}"

"d(e^xy)=-\dfrac{2e^x}{\cos x}dx"

Integrate


"\int d(e^xy)=-\int \dfrac{2e^x}{\cos x}dx"

"e^xy=-\int \dfrac{2e^x}{\cos x}dx"

"y=-e^{-x}\int \dfrac{2e^x}{\cos x}dx"